I required to design a PID controller that has overshoot less than 10% with minimise rise time, settling time, peak time and steady-state error.

The transfer function of the plant is shown below:

= 3.163 S^3 + 92.59 s^2 + 61.66 s

and the step response of the open loop system by using unit-step is shown below:

Step Response 40 35 30 25 Amplitude 20 15 10 5 0 0 100 200 500 600 700 300 400 Time (seconds)

Then I have designed my PID controller by referring to the example from Modern Control Engineering 5th Edition by Katsuhiko Ogata page 572 by using Ziegler Nichols 2nd Method.

I get Kcr = 5604.68 and Pcr = 0.8078. My PID controller transfer function is (339.5595*(s+4.9517)^2)/s. But the overshoot of my system exceeds 10%. How can I tune my PID controller so that I can get overshoot less than 10%?

Unit-Step Response 1.4 1.2 1 TED !!! 0.8 Amplitude 0.6 0.4 0.2 0 0 0.2 0.4 1 1.2 1.4 0.6 0.8 Time (seconds)
Answer #1

CODE :
clc close all s=tf(s); G = 3.163/(s^3+92.59*5^2+61.66*5); sisotool (G)]

REsult:

L SYSTEM VIEW ? model806 900 Editor - C:\Users\user\lab 3\model808.m +57 model803.m Xmodel804.m X model805.m clc close all 3

Control System Designer - IOTransfer_r2y: step Х CONTROL SYSTEM VIEW ? model806 C: Users user Editor - C:\Users\user\lab 3\mo

code:

clc close all s=tf(s); G = 3.163/(s^3+92.59*5^2+61.66*5); = = G_PID 7.4145+ (1+5.9*5+ (2.9*s)^2)/s; sys feedback (G_PID*G,

Х Figure 3 File Edit View Insert Tools Desktop Window Help 7 System: sys Peak amplitude: 1.06 ponse Overshoot (%): 5.98 At ti

Command Window New to MATLAB? See resources for Getting Started. ans = struct with fields: RiseTime: 0.9186 SettlingTime: 10.